Circular No. 7377

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SAX J1747.0–2853 C. B. Markwardt, University of Maryland and Goddard Space Flight

Center (GSFC); and F. E. Marshall and J. H. Swank, GSFC, report that RXTE PCA scans of the galactic bulge on Mar. 2.45 UT have detected a bright transient source near the galactic center. The position is consistent with a reappearance the bursting source SAX J1747.0–2853, discovered by the SAX Wide Field Camera in 1998 (IAUC 6846). The source was again detected in two follow-up pointed PCA observations and a subsequent galactic bulge scan. No significant pulsations or quasiperiodic oscillations were found. One x-ray burst was seen during the scan of Mar. 7.94. The approximate x-ray intensities in the band 2–10 keV were: Mar. 2.45, 42 mCrab; 6.37, 129; 7.94, 140; 8.38, 130. As the intensity of the source may still be increasing, further observations are encouraged.

SUPERNOVA 2000N IN MCG –02-34-054 S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for

Astrophysics, report that a low-S/N spectrum of SN 2000N (cf. IAUC 7354) was obtained by Jha on Mar. 8.4 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), showing a broad P-Cyg Hα feature, indicating that the supernova is of type II. Adopting the NASA/IPAC Extragalactic Database (NED) recession velocity for MCG –02-34-054 of 3990 km/s yields a supernova expansion velocity of ∼ 6000 km/s.

E. Cappellaro, S. Benetti, M. Turatto, A. Pastorello, S. Ortolani, and E. M. Corsini, Padua, report: “A spectrum (range 384–684 nm, resolution 1.0 nm) of SN 2000N, obtained on Mar. 9.2 UT with the Danish 1.54m telescope (+ DFOSC) at the European Southern Observatory, closely resembles that of a type-II supernova, ∼ 2 months past maximum. The expansion velocity deduced from Hα and Hβ absorption is ∼ 5000 km/s.”

COMET C/1999 E2 (SOHO) D. A. Biesecker, SM&A Corporation and Goddard Space Flight Cen-

ter, reports that M. Oates, Manchester, England, found another Kreutz sungrazing comet in archival SOHO LASCO C3 Web data. Astrometry and parabolic orbital elements appear on MPEC 2000-E39.

1999 UT  α_{2000}  δ_{2000}
Mar. 2.513  23^{h}19.^{m}1  − 8^{o}35′
2000 March 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green